The Baseball World Series championship is based on the best-of-seven games.

a. What is the probability that the World Series this year will last 4 games?

b. What is the probability that it will last 5 games?
Answer #1

Please don't hesitate to give a "thumbs up " for the answer, in case you're satisfied with it.

a. It will last 4 games, if by the time we play 4 games, majority of the 7 games is won.

So, 4 out 7 games are won by a team.

So, the probability of that happening is : (1/2)^4 , and since both teams can do this, we have the probability of this happening is , (1/2)^4 * 2 = (1/2)^3 = 1/8

b. P(it will last 5 games: the 4th game is won by a team in the 5th game:

4C3 (1/2)^3 (1/2)*(1/2) *2 = 4*1/2^4 = 1/4

So, basically, above, out of first 4 games, 3 were won by any of the 2 teams, and then the 5th game was won by that team.